if (A + B+ C+D) =pi, then ∑ cosA cosB - ∑sinA sinB =   

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Let A = B= C= D = pi/4

Value of required expression = 4 ( 0.5 ) - 4 ( 0.5 ) = 0

Regards